OpenStudy (anonymous):

its not letting me type into the post box, its a^2-11a+18=0

OpenStudy (anonymous):

ok to do these you have to find two numbers that add to -11 and multiply to 18

OpenStudy (anonymous):

a^2-9a-2a+18

OpenStudy (anonymous):

9 -2

OpenStudy (anonymous):

Now work out from here

OpenStudy (anonymous):

in this case it would be -2 and -9

OpenStudy (anonymous):

so put it in parenthesis (x-9)(x-2)=0

OpenStudy (anonymous):

Yeap a=9 and 2

OpenStudy (anonymous):

then set them against zero x-9=0 x=9 x-2=0 x=2 x={2,9}

OpenStudy (anonymous):

right it was a a={2,9}